Fermi National Laboratory
 

Operation Summary for March 29, 2010 - March 31, 2010

 

Monday March 29

The day shift began with Operations monitoring the Tevatron (TeV) store 7715 (86.83E30) and the Recycler stash (336.0E10), with the Antiproton Source stacking (21.19E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7715 at noon and put the TeV into HEP shot setup. Operations established store 7716 at 1 PM with an average initial luminosity (AIL) of 362E30. RF personnel repaired a MI RF station (MIRF12) and brought it back online.

During the evening shift, Operations disabled the MTest beam at 6 PM.

Tuesday March 30

The midnight shift began with Operations monitoring store 7716 (94.76E30) and the stash (269.7E10), with Pbar stacking (18.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 6:15 AM, the Linac LCW temperatures raised high enough to force Operations to turn off the low-energy and high-energy RF stations. Operators traced the LCW system problem to a failed pump at the Central Utility Building (CUB). The Duty Mechanic brought a backup pump online. Operators were bringing the Linac RF systems back on by the end of the shift.

During the day shift, Linac personnel brought LRF4 back online after recovering from the CUB pump failure. Operations restored beam to Pbar, NuMI, and MiniBooNE. Operations terminated store 7716 at 8:45 AM and put the TeV into shot setup. Operations established store 7717 at 10 AM with an AIL of 363E30. Operations established beam to MTest. MI personnel began beam studies at 3:40 PM.

During the evening shift, the MI beam studies concluded at 4:45 PM. Operations disabled the MTest beam at 8 PM.

Wednesday March 31

The midnight shift began with Operations monitoring store 7717 (80.09E30) and the stash (350.0E10), with Pbar stacking (23.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7717 at 2:50 AM and then put the TeV into shot setup. Operations established store 7718 at 4 AM with an AIL of 350E30.

The Plans for Wednesday

The plans for today are to stack, stash, and store. Preaccelerator personnel will remove and clean the H- Source.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/31/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory